OpenStudy (anonymous):

-3x6y4 x-3y-2 Which of the following is the simplest form of this expression? A.) -3x-2y-2 B.) -3x2y2 C.) -3x9y-2 D.) -3x9y6

OpenStudy (diyadiya):

Please use the equation tab to write down the equation, its confusing!

OpenStudy (anonymous):

Kay one sec...

OpenStudy (anonymous):

-3x^6y^4 / x^-3y^-2 in simplest form

OpenStudy (diyadiya):

\[\Large \frac{-3x^6y^4}{x^{-3}y^{-2}}\] is this your question?

OpenStudy (anonymous):

Yes yes yes!!!!!

OpenStudy (diyadiya):

Remember \[\Large x^{-m}= \frac{1}{x^m}\]Here m can be any number. and do you know what is ... ok let me give you an example \[\Large \frac{2}{ \frac{1}{9}} = ?\]

OpenStudy (anonymous):

I don't know i just need the equation in it's simplest form

OpenStudy (diyadiya):

Ok.. Anyways i won't be able to give away the final answer ,But i can make you understand. \[\Large x^{-m}= \frac{1}{x^m}\] For example : \[\Large x^{-3}= \frac{1}{x^3} ~~~~~ or ~~~~~ x^{-5} = \frac{1}{x^5}\] -------------------------------------------- \[\Large \frac{-3x^6y^4}{x^{-3}y^{-2}} =\Large \frac{-3x^6y^4}{\frac{1}{x^{3}} ~~ \frac{1}{y^{2}}}\] Did you understand this

OpenStudy (diyadiya):

and you should also know that \[\Large x^m \times x^n = x^{m+n}\]